OpenStudy (anonymous):

In how many ways can a mail carriers place three letters in five mailboxes if: a) each box may receive more than one letter? b) each mailbox may receive no more than one letter?

OpenStudy (anonymous):

Evaluate by De Moivre’s Theorem and express the result in rectangular form.\[\left\{ 3\left( \frac{ \sqrt{2} }{ 2 }-j \frac{ \sqrt{2} }{ 2 } \right)^{\frac{ 1 }{ 3 }} \right\}^{\frac{ 1 }{ 2 }}\]